1995 AISB Workshop Series Automated Reasoning: Bridging the Gap between Theory and Practice List of Workshop Participants

Jonas Barklund Uppsala University Dept. of Computing Science Box 411, S-751 05 Uppsala Sweden Email: jonas@csd.uu.se

Brandon Bennett Division of Artificial Intelligence School of Computer Studies University of Leeds Leeds, LS2 9TJ United Kingdom Email: brandon@scs.leeds.ac.uk

Narjes Berregeb CRIN & INRIA-Lorraine BP 239 54506 Vandoeuvre-les-Nancy France Email: Narjes.Berregeb@fr.loria

Marc Bezem Dept. Philosophy University of Utrecht Heidelberglaan 8 3584 CS Utrecht The Netherlands Email: Marc.Bezem@nl.ruu.phil

Wolfgang Bibel Technische Hochschule Darmstadt FB 20 / Intellektik Alexanderstr. 10 D-64283 Darmstadt Germany Email: bibel@de.th-darmstadt.informatik.intellektik James E Borrett Dept. of Computer Science University of Essex, Wivenhoe Park, Colchester, Essex, CO4 3SQ United Kingdom Email: jborrett@essex.ac.uk

Alan Bundy

Dept. Artificial Intelligence University of Edinburgh 80 South Bridge Edinburgh, EH1 1HN United Kingdom Email: A.Bundy@ed.ac.uk

Anthony Cohn

Division of Artificial Intelligence School of Computer Studies University of Leeds Leeds, LS2 9TJ United Kingdom Email: agc@scs.leeds.ac.uk

Maria R. Cravo

Dept. de Mecanica - Instituto Superior Tecnico Av. Rovisco Pais 1000 Lisboa Portugal Email: MRCravo@poirot.ist.utl.pt Paul Curzon University of Cambridge Computer Laboratory New Museums Site Pembroke Street Cambridge United Kingdom Email: Paul.Curzon@cl.cam.ac.uk

Andrew Davenport Department of Computer Science University of Essex Colchester Essex, CO4 3SQ United Kingdom Email: daveat@essex.ac.uk

David Duffy Dept. of Computer Science University of York York, YO1 5DD United Kingdom Email: dad@minster.york.ac.uk

Edmund Dumbill Department of Computer Science University of York York, YO1 5DD United Kingdom Email: ejad-a@minster.york.ac.uk

Saturnino F Luz Filho University of Edinburgh - CCS 2 Buccleuch Place Edinburgh, EH8 9LW United Kingdom Email: luz@cogsci.ed.ac.uk

Michael Fisher Dept. of Computing Manchester Metropolitan University Chester Street Manchester, M1 5GD United Kingdom Email: M.Fisher@doc.mmu.ac.uk Enrico Franconi Knowledge Representation and Reasoning group IRST, Istituto per la Ricerca Scientifica e Tecnologica I-38050 Povo TN, ITALY Email: franconi@irst.it

Alan Frisch

Department of Computer Science

University of York York, YO1 5DD United Kingdom Email: frisch@minster.york.ac.uk

Ian P. Gent

Dept. of Computer Science University of Strathclyde Glasgow, G1 1XH United Kingdom Email:ipg@cs.strath.ac.uk

Stuart A. Grant

Division of Artificial Intelligence School of Computer Studies University of Leeds Leeds, LS2 9TJ United Kingdom Email: stuartg@scs.leeds.ac.uk

Alan G Hamilton

Dept. of Computing Science

University of Stirling Stirling, FK9 4LA United Kingdom Email: a.g.hamilton@stir.cs

Pat Hill

School of Computer Studies

University of Leeds Leeds, LS2 9TJ United Kingdom Email: hill@scs.leeds.ac.uk Dieter Hutter DFKI Saarbrucken Stuhlsatzenhausweg 3 D - 66123 Saarbruecken Germany Email: hutter@de.uni-sb.dfki

Francisco Ibanez School of Computer Studies University of Leeds Leeds, LS2 9JT United Kingdom Email: pancho@scs.leeds.ac.uk

Andrew Ireland Dept. Artificial Intelligence University of Edinburgh 80 South Bridge Edinburgh, EH1 1HN United Kingdom Email: a.ireland@ed.ac.uk

Robert Johnson Dept. of Computing Manchester Metropolitan University Chester Street Manchester, M1 5GD United Kingdom Email: R.Johnson@doc.mmu.ac.uk

Raul Lopez Division of Artificial Intelligence University of Leeds Leeds, LS2 9TJ United Kingdom Email: raul@scs.leeds.ac.uk

Zhaohui Luo Dept. of Computer Science University of Durham South Road Durham, DH1 3LE United Kingdom Email: zhaohui.luo@durham.ac.uk Craig MacNish Dept. of Computer Science University of York York, YO1 5DD United Kingdom Email: craig@minster.york.ac.uk

Simin Nadjm-Tehrani Dept. of Computer and Information Science, Linkoeping University, S-581 83 Linkoeping, Sweden. Email: snt@se.liu.ida

Maurice Naftalin Dept. Artificial Intelligence University of Edinburgh 80 South Bridge Edinburgh, EH1 1HN United Kingdom Email: mauricen@ed.aisb

Miles Osborne Dept. of Computer Science University of York York, YO1 5DD United Kingdom Email: miles@minster.york.ac.uk

S. Poria LOLITA reasoning group, Laboratory for Natural Language Engineering, Department of Computer Science, University of Durham Science Site, Stockton Road, Durham, DH1 3LE United Kingdom Email: Sanjay.Poria@durham.ac.uk

Patrick Prosser Department of Computer Science University of Strathclyde Glasgow, G1 1XH United Kingdom Email: pat@cs.strath.ac.uk Giovanni Semeraro Dipartimento di Informatica Universita’ di Bari Italy Email: semeraro@it.uniba.plotkin

S. Shiu LOLITA reasoning group, Laboratory for Natural Language Engineering, Department of Computer Science, University of Durham Science Site, Stockton Road, Durham, DH1 3LE United Kingdom Email: S.K.Shiu@durham.ac.uk

Barbara Smith Division of Artificial Intelligence School of Computer Studies University of Leeds Leeds, LS2 9TJ United Kingdom Email: bms@scs.leeds.ac.uk

Werner Stephan DFKI Saarbrucken Stuhlsatzenhausweg 3 D - 66123 Saarbruecken Germany Email: stephan@de.uni-sb.dfki

Frieder Stolzenburg Universitaet Koblenz-Landau Rheinau 1 D-56075 Koblenz Germany Email: stolzen@informatik.uni-koblenz.de

H.C.M de Swart Dept. of Philosophy Tilburg University PO Box 90153 5000 LE Tilburg The Netherlands Email: swart@kub.nl Mark Tarver Division of Artificial Intelligence University of Leeds Leeds, LS2 9TJ United Kingdom Email: mark@leeds.scs

Judith Underwood

Dept. of Computer Science University of Edinburgh King’s Buildings Edinburgh, EH9 3JZ United Kingdom Email: jlu@dcs.ed.ac.uk

Laurent Vigneron

Department of Computer Science SUNY at Stony Brook Stony Brook, NY 11794 USA Email: lavig@cs.sunysb.edu

Toby Walsh

Mechanized Reasoning Group IRST Trento & DIST, University of Genoa Italy Email: toby@irst.it